1. You must analyze the cash flows of two projects, S and L. Project S: CF0 = -1500; CF1 = 800; CF2 = 700; CF3 = 100; CF4 = 600 Project L: CF0 = -1500; CF1 = 200; CF2 = 600; CF3 = 900; CF4 = 700 Given a required rate of return of 10%, what is the IRR of the better project? (Note: the better project may not be the one with the higher IRR)
Answer #1

Better project is the one that has higher NPV

Project S:

NPV = -1500 + 800 / (1 + 0.1)1 + 700 / (1 + 0.1)2 + 100 / (1 + 0.1)3 + 600 / (1 + 0.1)4

NPV = $290.73

Project L:

NPV = -1500 + 200 / (1 + 0.1)1 + 600 / (1 + 0.1)2 + 900 / (1 + 0.1)3 + 700 / (1 + 0.1)4

NPV = $331.98

Better project is project L as it has a higher NPV.

IRR is the rate of return that makes NPV equal to 0.

NPV = -1500 + 200 / (1 + R)1 + 600 / (1 + R)2 + 900 / (1 + R)3 + 700 / (1 + R)4

Using trial and error method, let's try R as 18.27%

NPV = -1500 + 200 / (1 + 0.1827)1 + 600 / (1 + 0.1827)2 + 900 / (1 + 0.1827)3 + 700 / (1 + 0.1827)4

NPV = 0

Therefore, IRR of project L is 18.27%